# JointDiff: Bridging Continuous and Discrete in Multi-Agent Trajectory Generation

### v1

@@ D.5.2 Controllable Generation
D.5.3 Failure Cases
A key limitation observed in our results is the occasional inconsistency between the generated trajectories and the text guidance. This can be traced to the constrained size of our training datasets, $\approx 10$k pairs for NFL and $\approx 2$k (augmented to 4k) for BundesLiga. The limited data variety hinders the model’s ability to robustly encode the wide range of possibilities described in natural language. Refer to Fig.14 to see some failure examples from the same scenarios depicted in Fig. 13.
@@ Appendix E Limitations and Future Work
Our model’s architecture requires events to share the same spatio-temporal structure as the trajectory data, i.e., to allow for simple concatenation at the input. This limits its application to event streams that are naturally structured this way and cannot directly handle unstructured data, such as sparse temporal point processes (TPPs). A key direction for future work is to develop methods to integrate these more complex event types.

### v2

@@ D.5.1 Completion Generation
D.5.2 Controllable Generation
D.5.3 Failure Cases
A key limitation observed in our results is the occasional inconsistency between the generated trajectories and the text-guidance. This can be traced to the constrained size of our training datasets, $\approx 10$k pairs for NFL and $\approx 2$k (augmented to 4k) for Bundesliga. The limited data variety hinders the model’s ability to robustly encode the wide range of possibilities described in natural language. Refer to Fig.14 to see some failure examples from the same scenarios depicted in Fig. 13.
@@ Appendix E Limitations and Future Work
Our model’s architecture requires events to share the same spatio-temporal structure as the trajectory data, i.e., to allow for simple concatenation at the input. This limits its application to event streams that are naturally structured this way and cannot directly handle unstructured data, such as sparse temporal point processes. A key direction for future work is to develop methods to integrate these more complex event types.